9. Explain as much as you can! I alreadyknow the answer, but need to know why! Use written words as well.sp (9) Calculate the pH of a buffer solution that
Answer #1

Acidic buffer is a solution of a weak acid and its salt. The Henderson-Hasselbalch approximation can be used to calculate the pH of a buffer solution. The balanced equation for an acidic buffer is:

HAH++A

The strength of a weak acid is usually represented as an equilibrium constant. The acid-dissociation equilibrium constant

Ka = [H+][A-]/[HA]

or, -logKa = -log[H+] -log[A-]/[HA]

or, pKa = pH - log[A-]/[HA]

or, pH = pKa + log [A-]/[HA]

HA is a weak acid. By definition, HA does not dissociate completely and we can write,

[HA] = [HA]initial

[A-] = [A-]initial

So, pH = pKa + log [A-]/[HA] = pH = pKa + log [A-]initial/[HA]initial

= pKa + log[salt]/[acid]

This equation is known as Henderson-Hasselbalch equation.

According to the queation,

pKa of acidic acid = -logKa = -log (1.77*10-5) = 4.75

We need to calculate the concentrations of the salt(sodium acetate) and the acid first.

Molar mass of sodium acetate = 82 g/mol

So, 0.82 g sodium acetate = 0.82/82 mole = 0.01 mol

100 mL solution contains 0.01 moles of sodium acetate. So [sodium acetate]= 0.01*1000/100 = 0.1 M

[acid] = 0.010 * 1000/100 = 0.1 M

So, pH = pKa + log[salt]/[acid]

= 4.75 + log 0.1/0.1

= 4.75